# SqueezeAttention: 2D Management of KV-Cache in LLM Inference via Layer-wise Optimal Budget

### v1

@@ Abstract
Optimizing the Key-Value (KV) cache of the Large Language Model (LLM) has been considered critical to saving the cost of inference. Most of the existing KV-cache compression algorithms attempted to sparsify the sequence of tokens by taking advantage of the different importance of tokens. In this work, we found that by identifying the importance of attention layers, we could optimize the KV-cache jointly from two dimensions. Based on our observations regarding layer-wise importance in inference, we propose SqueezeAttention to precisely optimize the allocation of KV-cache budget among layers on-the-fly and then incorporate three representative token sparsification algorithms to compress the KV-cache for each layer with its very own budget. By optimizing the KV-cache from both sequence’s and layer’s dimensions, SqueezeAttention achieves around 30% to 70% of the memory reductions and up to 2.2 $\times$ of throughput improvements in a wide range of LLMs and benchmarks. The code is available at https://github.com/hetailang/SqueezeAttention.
SqueezeAttention: 2D Management of KV-Cache in LLM Inference via Layer-wise Optimal Budget
Zihao Wang Lanzhou University wangzihao2020@lzu.edu.cn                        Shaoduo Gan thanks: Corresponding author. Geoming AI ganshaoduo@gmail.com
Figure 1: Demonstrations of KV-cache policies in inference from the view of the sequence and attention layer. Full Cache (leftmost column) simply stores the KV embeddings for all the tokens in all the layers. Intra-layer compression algorithms (middle column) drop tokens in the sequence’s dimension, where each layer has the same cache budget. SqueezeAttention (rightmost column) further compresses the KV-cache by adaptively re-allocating the cache budgets in the layer’s dimension.
@@ 1 Introduction
For a decoder-only autoregressive model, which is the most widely adopted LLM architecture, the inefficiencies in inference mainly come from the fact that the model can only generate tokens one by one, and sampling each token requires attending to all previous tokens. In practice, the intermediate key-value embeddings of each layer have been cached incrementally in each iteration to avoid frequent recomputations. Since the KV-cache increases linearly with the number of attention layers, context length and batch size, it often ends up being multiple times larger than the model itself Sheng et al. (2023), and therefore, dominating the I/O cost of inference. Recently, optimizing the KV-cache has been broadly considered a critical approach to boost the efficiency of inference. From the perspective of context length, many well-studied algorithms are trying to identify the most "valuable" tokens in the sequence and evict the unimportant ones to reduce the KV-cache and attention complexity, such as Sliding Window Attention Beltagy et al. (2020), Heavy-Hitter (H2O) Zhang et al. (2024), StreamingLLM Xiao et al. (2023), Scissorhands Liu et al. (2024), FastGen Ge et al. (2023) and so on. From the perspective of batching, many studies aim to explore how to efficiently manage the memory of KV-cache on a batch basis with different sequence lengths Zheng et al. (2023); Kwon et al. (2023). However, the opportunities in the dimension of attention layers have barely been touched by most, if not all, of the existing methods. In other words, all the attention layers have always been treated equally by those KV caching strategies. Therefore, in this paper we ask:
@@
Do all the attention layers necessarily have to store the same amount of KV-cache? If not, how can we precisely reallocate the cache budget for each layer such that we can further reduce the KV-cache in total?
@@
To answer these questions, we need to take a closer look at the behaviors of different attention layers during inference. Some inspiring clues could be found in a few related studies. Early-exiting LLM Del Corro et al. (2023), as a widely-adopted inference method, shows that after going through a certain number of attention layers, the hidden representations are likely to reach saturation, and therefore, the forward computing can exit early without finishing the entire network and still get a reasonable prediction. Besides, a very recent work called FastGen Ge et al. (2023) found that attention layers in different positions have quite different preferences regarding KV caching policies. For example, attention heads in the first and last layers prefer caching all the tokens in the sequence, whereas, some middle layers prefer caching only frequent and special tokens. These insights indicate that attention layers do have quite different degrees of importance regarding inference, encouraging us to make a hypothesis that by taking advantage of the layer importance we could further "squeeze" the KV-cache that has already been optimized by those existing intra-layer algorithms, from the layer’s dimension, and eventually, achieve even better efficiencies.
@@
Based on this simple yet effective indicator, we propose SqueezeAttention, a 2D KV-cache compression algorithm that prunes KV-cache from not only the sequence’s dimension but also the layer’s dimension. Since the layer importance is highly dependent on the model and task, SqueezeAttention categorizes all the layers into groups on the fly by clustering their cosine similarities measured during the prompt prefilling phase. Given an intra-layer KV-cache compression policy (like Sliding Window Beltagy et al. (2020) or H2O Zhang et al. (2024)), and a unified cache budget (like 4096 tokens or 20% of prompt length), SqueezeAttention automatically reallocates the cache budgets among groups of layers such that the important layers could cache more tokens to stabilize the model accuracy and the unimportant layers could drop more unnecessary tokens to save the I/O cost. What’s even better is that SqueezeAttention is orthogonal to all those intra-layer KV-cache compression algorithms, so it can be smoothly combined with any of them. Figure 1 demonstrates how the SqueezeAttention works jointly with two representative KV compression algorithms, i.e., H2O Zhang et al. (2024) and Sliding Window Attention Beltagy et al. (2020). More details about the SqueezeAttention algorithm can be found in Section 4.
To the best of our knowledge, SqueezeAttention is the first algorithm considering the KV-cache budget in a layer-wise way, making it a valuable addition to all those sequence-wise sparsification algorithms for inference. In our experiments, we integrate SqueezeAttention into 7 popular LLM models, i.e., Llama2-7B, Mistral-7B, Falcon-7B, OPT-6.7B, GPT-Neox-20B, Mixtral-8$\times$7B, and Llama2-70B, combining with 3 representative intra-layer KV-cache compression algorithms, i.e., Heavy-Hitter Oracle (H2O), Sliding Window Attention and StreamingLLM. The results show that SqueezeAttention can achieve better model performance with even lower cache budgets than all three algorithms under a wide range of models and tasks, which lead to approximately 30% to 70% of the memory savings and up to 2.2 $\times$ of throughput improvements for inference.
@@ 2.2 Existing KV-cache Optimizations
Sparsifying the context sequence is an effective way to break the linear relationship between the context length and the KV-cache Del Corro et al. (2023); Zhang et al. (2024); Anagnostidis et al. (2023); Sukhbaatar et al. (2019); Rae and Razavi (2020). The general intention of these algorithms is to find out the unimportant tokens in the sequence and drop the KV-cache of these tokens. For example, Sliding Window Attention Beltagy et al. (2020) only caches a certain number of the most recent tokens and drops the rest. StreamingLLM Xiao et al. (2023) found that in addition to the recent tokens, tokens at the beginning of the sequence are also crucial to the output. Heavy-Hitter Zhang et al. (2024) and Scissorhands Liu et al. (2024) rank the importance of tokens by comparing their attention scores. As mentioned above, these algorithms treat all the attention layers equally, and therefore, have a fixed KV-cache budget for each layer.
@@
Last but not least, how to relax the linear relationship between the KV-cache and the layers remains largely unexplored compared with the other two dimensions. FastGen Ge et al. (2023) is a very recent work that found layers in different positions may have different preferences on the KV-cache policies. It then proposed an algorithm to choose the best cache policy for each attention head during the inference. Note that FastGen aims to find out the optimal cache policy for attention heads, but all the heads sharing the same policy still simply have the same cache budget, like 30% of the most recent tokens, which makes SqueezeAttention still orthogonal to it.
@@ 3 Observations
$=\frac{\sum\limits_{i=1}^{n}A_{i}\cdot B_{i}}{\sqrt{\sum\limits_{i=1}^{n}A_{i}^{2}}\cdot\sqrt{\sum\limits_{i=1}^{n}B_{i}^{2}}}$ (6)
@@
Figure 3: Performance of SqueezeAttention, baseline algorithms, and Full Cache under different cache budgets.
@@ 4 Algorithm
In the sequence’s dimension, there are various cache eviction policies that we could directly combine with, like Least Recently Used methods (Sliding Window, StreamingLLM), Least Frequently Used methods (Scissorhands, H2O), and so on. We denote a policy that compresses the KV-cache in sequence’s dimension by $\boldsymbol{C}_{seq}$, and its cache budget by $b_{init}$. Note that all the layers have the same cache budget by default, just like the assumptions made by all these intra-layer KV compression algorithms.
In the layer’s dimension, SqueezeAttention firstly tracks the layer importance with the given prompt in the prefilling phase by collecting the cosine similarities of each layer whenever the self-attention is conducted. At the end of prefilling, each layer ends up with a set of cosine similarities, each of which corresponds to a token that has flowed through this layer. Then we use the averaged value over prompt tokens to represent the layer-wise importance of this layer regarding this task. By clustering the layers into groups based on the layer-wise importance with KMeans, we reallocate the $b_{init}$ for each layer in a way that more budgets are assigned to the more "important" layer groups. Since layers have different cache budgets, in the decoding phase, $\boldsymbol{C}_{seq}$ works separately with each layer’s very own budgets. The detailed process is described in Algorithm 1.
@@
1:$prompt$: input sequence; $\boldsymbol{C}_{seq}$: a KV-cache compressor in sequence dimension; $b_{init}$: the initial cache budget of each layer; $n_{layer}$: number of attention layers; $p$: hyperparameter ($10<p<$ 1); $K^{(i)}$: KV-cache of the $i$-th layer;
@@
15:for $o\leftarrow 1$ to $\texttt{len}(output)$ donormal-$\triangleright$ Decoding output tokens one by one
@@
18:              $K^{(i)}=\boldsymbol{C}_{seq}(K^{(i)},b^{(i)})$triangleright Compress the KV-cache of this layer by its own budget
@@
Discussions. SqueezeAttention involves a hyperparameter $p$ to control the percentage of initial budgets that could be removed from the "unimportant" layers. The bigger the $p$ is, the more budgets will be reassigned. In experiments, we found 0.1-0.2 is a reasonable choice range in most cases. Empirically, we cluster layers into 3 groups. Group 1 (smallest similarity) usually stands for a few special layers, like the first and last layers, whose similarities are way lower than other layers. Group 3 (biggest similarity, typically between 0.95 to 0.99) represents the layers that barely make any difference to the embeddings after the self-attention. Group 2 is in the middle. To preserve the model accuracy, we only reduce cache budgets from Group 3, which accounts for around 50% to 70% of total layers.
@@
Table 2: Performance among Full cache, SqueezeAttention and Baseline Algorithms.
Model Size Dataset Baseline Algorithm Performance / Used KV Budget
Full Cache SqueezeAttention Basic Algorithm
@@
(a) LLama2-70B
(b) Mistral-7B
(c) GPT-NeoX-20B
Figure 4: Comparisons of per-token decoding memory usage among the Full cache, SqueezeAttention, and baseline algorithms, all methods achieve similar accuracies shown in Table 2.
5 Experiments
In this section, we evaluate the SqueezeAttention on a wide range of LLM models and datasets. We choose 3 state-of-the-art sequence-wise compression algorithms as baselines, and show how our algorithm works jointly with them to achieve even better performance regarding the model accuracy and hardware efficiency.
5.1 Experiment Setup
Baselines. 3 sequence-wise sparsification algorithms are chosen as the baselines to integrate into SqueezeAttention:
  • Heavy-Hitter (H2O): Identify the important tokens within the sequence by comparing the accumulated attention score of each token.
  • Sliding Window Attention: A "Local" strategy that only caches the most recent tokens’ KV embeddings. This method works especially well with Mistral and Mixtral.
  • StreamingLLM: In addition to the most recent tokens, StreamingLLM always caches the first $n$ tokens in the sequence, as they are identified as "sink tokens". We take $n=4$ as recommended by the paper.
5.2 End-to-End Result
Figure 3 demonstrates the comparison results of SqueezeAttention and other 3 baseline algorithms over all 7 models and 5 datasets. Full Cache (dashed line) means all tokens’ KV embeddings are fully cached during the inference, therefore, it represents a relatively good model accuracy. The blue and orange curves in each subfigure illustrate how the model accuracy changes with the KV-cache budgets ranging from 10% to 100% of the total sequence length. For each model, we choose the best sequence-wise compression algorithm to represent the best case of intra-layer optimization, and then use SqueezeAttention on top of the best case. As shown in the figure, SqueezeAttention consistently manages to improve the model accuracies under various KV-cache budgets by reallocating the cache budgets among layers. In other words, SqueezeAttention can achieve similar inference accuracies with much less KV-cache in total.
Table 3: Generation throughput (token / s) on eight A100 GPUs of Mistral-7B and LLama2-70B with SqueezeAttention and Full Cache. To maintain the model accuracy, SqueezeAttention uses 20% of the cache budget for Mistral-7B and 30% of the cache budget for LLama2-70B. "OOM" means out-of-memory.

### v2

@@ Abstract
Optimizing the Key-Value (KV) cache of the Large Language Model (LLM) has been considered critical to saving the cost of inference. Most of the existing KV-cache compression algorithms attempted to sparsify the sequence of tokens by taking advantage of the different importance of tokens. However, most of these methods treat all layers equally, allocating the same KV budget to each layer. This approach is suboptimal, as some layers may be less sensitive to input tokens yet still receive the same budget as others. In this work, we found that by identifying the importance of attention layers, we could optimize the KV-cache jointly from two dimensions, i.e., sequence-wise and layer-wise. Based on our observations regarding layer-wise importance in inference, we propose SqueezeAttention to precisely optimize the allocation of KV-cache budget among layers on-the-fly and then incorporate three representative sequence-wise algorithms to compress the KV-cache for each layer with its very own budget. Specifically, we first measure each layer’s importance by calculating the cosine similarity of the input prompt differences before and after the self-attention layers. Based on this similarity, we then categorize the layers into two groups and adjust their KV budgets accordingly. By optimizing the KV-cache from both sequence’s and layer’s dimensions, SqueezeAttention achieves around 30% to 70% of the memory reductions and up to 2.2 $\times$ of throughput improvements in a wide range of LLMs and benchmarks. The code is available at https://github.com/hetailang/SqueezeAttention.
**footnotetext: This work is done during the internship at Peking University and Geoming AI.$\dagger$\$$\dagger$\$footnotetext: Corresponding author.
@@ 1 Introduction
For a decoder-only autoregressive model, which is the most widely adopted LLM architecture, the inefficiencies in inference mainly come from the fact that the model can only generate tokens one by one, and sampling each token requires attending to all previous tokens. In practice, the intermediate key-value embeddings of each layer have been cached incrementally in each iteration to avoid frequent recomputations. Since the KV-cache increases linearly with the number of attention layers, context length and batch size, it often ends up being multiple times larger than the model itself Sheng et al. (2023), and therefore, dominating the I/O cost of inference. Recently, optimizing the KV-cache has been broadly considered a critical approach to boost the efficiency of inference. From the perspective of context length, many well-studied algorithms are trying to identify the most ”valuable” tokens in the sequence and evict the unimportant ones to reduce the KV-cache and attention complexity, such as Sliding Window Attention Beltagy et al. (2020), Heavy-Hitter (H2O) Zhang et al. (2024), StreamingLLM Xiao et al. (2023), Scissorhands Liu et al. (2024), FastGen Ge et al. (2023) and so on. From the perspective of batching, many studies aim to explore how to efficiently manage the memory of KV-cache on a batch basis with different sequence lengths Zheng et al. (2023); Kwon et al. (2023). However, the opportunities in the dimension of attention layers have barely been touched by most, if not all, of the existing methods. In other words, all the attention layers have always been treated equally by those KV caching strategies. Therefore, in this paper we ask:
Figure 1: Demonstrations of KV-cache policies in inference from the view of the sequence and attention layer. Full Cache (leftmost column) simply stores the KV embeddings for all the tokens in all the layers. Sequence-wise compression algorithms (middle column) drop tokens in the sequence’s dimension, where each layer has the same cache budget. SqueezeAttention (rightmost column) further compresses the KV-cache by adaptively re-allocating the cache budgets in the layer’s dimension.
@@
Do all the attention layers that share the same KV caching strategy have to cache the same amount of tokens? If not, how can we precisely allocate the cache budget for each layer such that we can further reduce the KV-cache on top of sequence-wise compressions?
@@
To answer these questions, we need to take a closer look at the behaviors of different attention layers during inference. Some inspiring clues could be found in a few existing studies. Early-exiting LLM Del Corro et al. (2023), as a widely-adopted inference method, shows that after going through a certain number of attention layers, the hidden representations are likely to reach saturation, and therefore, the forward computing can exit early without finishing the entire network and still get a reasonable prediction. Besides, a very recent work called FastGen Ge et al. (2023) found that attention layers in different positions have different optimal KV caching strategies. For example, attention layers in the very early part of the model should simply cache all the tokens in the sequence, whereas, some middle layers should apply cache eviction strategies based on the token locality or frequency. Although FastGen could select the optimal eviction strategy for each layer, it is still unclear how to find the optimal cache budgets, instead of a pre-defined, unified hyperparameter, for layers that share the same strategy.
Given the fact that attention layers do have different degrees of importance regarding inference, we can make a reasonable hypothesis that by taking advantage of the layer importance, we could further ”squeeze” the amount of KV-cache that has already been compressed by those sequence-wise eviction algorithms, and eventually, achieve even better efficiencies.
@@
Based on this simple yet effective indicator, we propose SqueezeAttention, a 2D KV-cache compression algorithm that prunes KV-cache from not only the sequence’s dimension but also the layer’s dimension. Since the layer importance is highly dependent on the model and task, SqueezeAttention categorizes all the layers into groups on the fly by clustering their cosine similarities measured during the prompt prefilling phase. Given a sequence-wise KV-cache eviction policy (like Sliding Window Beltagy et al. (2020) or H2O Zhang et al. (2024)), and a unified cache budget (like 4096 tokens or 20% of prompt length), SqueezeAttention automatically reallocates the cache budgets among groups of layers such that the important layers could cache more tokens to stabilize the model accuracy and the unimportant layers could drop more tokens to save the I/O cost. What’s even better is that SqueezeAttention is orthogonal to all those sequence-wise KV-cache compression algorithms, so it can be smoothly combined with any of them. Figure 1 demonstrates how the SqueezeAttention works jointly with two representative sequence-wise KV-cache eviction algorithms, i.e., H2O Zhang et al. (2024) and Sliding Window Attention Beltagy et al. (2020). More details about the SqueezeAttention algorithm can be found in Section 4.
To the best of our knowledge, SqueezeAttention is the first algorithm considering the KV-cache budget in a layer-wise way, making it a valuable addition to all those sequence-wise compression algorithms for inference. In our experiments, we integrate SqueezeAttention into 7 popular LLM models ranging from 7B to 70B, i.e., Llama2-7B, Mistral-7B, Falcon-7B, OPT-6.7B, GPT-Neox-20B, Mixtral-8$\times$7B, and Llama2-70B, combining with 3 representative sequence-wise KV-cache compression algorithms, i.e., Heavy-Hitter Oracle (H2O), Sliding Window Attention and StreamingLLM. The results show that SqueezeAttention can achieve better model performance with even lower cache budgets than all three algorithms under a wide range of models and tasks, which lead to approximately 30% to 70% of the memory savings and up to 2.2 $\times$ of throughput improvements for inference.
@@ 2.2 Existing KV-cache Optimizations
Sparsifying the context sequence is an effective way to break the linear relationship between the context length and the KV-cache Del Corro et al. (2023); Zhang et al. (2024); Anagnostidis et al. (2023); Sukhbaatar et al. (2019); Rae & Razavi (2020). The general intention of these algorithms is to find out the unimportant tokens in the sequence and drop the KV-cache of these tokens. For example, Sliding Window Attention Beltagy et al. (2020) only caches a certain number of the most recent tokens and drops the rest. StreamingLLM Xiao et al. (2023) found that in addition to the recent tokens, tokens at the beginning of the sequence are also crucial to the output. Heavy-Hitter Zhang et al. (2024) and Scissorhands Liu et al. (2024) rank the importance of tokens by comparing their attention scores. As mentioned above, these algorithms treat all the attention layers equally, and therefore, have a fixed KV-cache budget for each layer.
@@
Last but not least, how to relax the linear relationship between the KV-cache and the layers remains largely unexplored compared with the other two dimensions. FastGen Ge et al. (2023) is a very recent work that found layers in different positions may have different optimal sequence-wise KV-cache eviction strategies. It then proposed an algorithm to choose the best eviction strategy from 1) Locality strategy (like Sliding Window), 2) Special Tokens strategy (like StreamingLLM), 3) Local and Frequency strategy (like Scissorhands, H2O) and so on for each attention head during the inference. However, for all the attention heads that have been assigned the same eviction strategy, FastGen simply gives them a unified pre-defined cache budget, like 30% of the sequence length. Therefore, how to adaptively allocate the KV-cache budget to layers with the same sequence-wise token eviction strategy is still largely unclear.
@@ 4 Algorithm
4.1 SqueezeAttention
In the sequence’s dimension, there are various cache eviction policies that we could directly combine with, like Least Recently Used methods (Sliding Window, StreamingLLM), Least Frequently Used methods (Scissorhands, H2O), and so on. We denote a policy that compresses the KV-cache in sequence’s dimension by $\bm{C}_{seq}$, and its cache budget by $b_{init}$. Note that all the layers have the same cache budget by default, just like the assumptions made by all these sequence-wise KV-cache compression algorithms.
In the layer’s dimension, SqueezeAttention firstly tracks the layer importance with the given prompt in the prefilling phase by collecting the cosine similarities of each layer whenever the self-attention is conducted. At the end of prefilling, each layer ends up with a set of cosine similarities, each of which corresponds to a token that has flowed through this layer. Then we use the averaged value over prompt tokens to represent the layer-wise importance of this layer regarding this task. By clustering the layers into groups based on the layer-wise importance with KMeans, we reallocate the $b_{init}$ for each layer in a way that more budgets are assigned to the more ”important” layer groups. Since layers have different cache budgets, in the decoding phase, $\bm{C}_{seq}$ works separately with each layer’s very own budgets. The detailed process is described in Algorithm 1.
@@ 4.1 SqueezeAttention
1:$prompt$: input sequence; $\bm{C}_{seq}$: a KV-cache compressor in sequence dimension; $b_{init}$: the initial cache budget of each layer; $n_{layer}$: number of attention layers; $p$: hyperparameter ($10<p<$ 1); $K^{(i)}$: KV-cache of the $i$-th layer;
@@
15:for $o\leftarrow 1$ to $\texttt{len}(output)$ $\triangleright$ Decoding output tokens one by one
@@
18:              $K^{(i)}=\bm{C}_{seq}(K^{(i)},b^{(i)})$triangleright Compress the KV-cache of this layer by its own budget
@@
4.2 Discussions
4.2.1 How to decide the value of $p$
SqueezeAttention involves a hyperparameter $p$ to control the percentage of initial budgets that could be removed from the ”unimportant” layers. The smaller the $p$ is, the more budgets will be reassigned. In experiments, we found 0.3-0.4 is a reasonable choice range in most cases. To precisely understand the impact of $p$, we have conducted extra experiments to demonstrate how the model accuracy changs with the value of $p$, please refer to A.2 for more details.
4.2.2 Why clustering into three groups?
Based on the observations of 7 models we have tried, we found they all have a typical pattern (3 groups) with respect to the layer importance. Specifically, Group 1 consists of a few special layers (always the first and last few layers) which can be seen as an analogy of special tokens that should never be evicted. Then Group 2 and Group 3 do not have a fixed borderline with each other, but we found that Group 3 makes obviously less impact on the embeddings, which can be seen as an analogy of ”frequent” and ”unfrequent” tokens in the sequence-wise methods. Therefore, our 3-group policy could be rephrased as: ”we firstly identify and prioritize the special layers (Group 1), then classify the rest layers into two groups: important (Group 2) and unimportant (Group 3), then reallocate the cache budget based on the clustering. Even if we cluster the layers into more than 3 groups, we just break down Group 2 and Group 3 into many small groups, but they eventually need to be reduced into two classes again, that is, either reducing the budget or increasing the budget. To preserve the model accuracy, we only reduce cache budgets from Group 3, which accounts for around 50% to 70% of total layers.
@@ 4.2.2 Why clustering into three groups?
5 Experiments
5.1 Experiment Setup
Figure 3: Performance of SqueezeAttention, best baselines, and Full Cache under different cache budgets.
Baselines. 3 sequence-wise sparsification algorithms are chosen as the baselines to integrate into SqueezeAttention and we assign each algorithm to the model that works beat, which we call the beat baseline algorithms:
  • Heavy-Hitter (H2O) Zhang et al. (2024): Identify the important tokens within the sequence by comparing the accumulated attention score of each token.
  • Sliding Window Attention Beltagy et al. (2020): A ”Local” strategy that only caches the most recent tokens’ KV embeddings. This method works especially well with Mistral and Mixtral.
  • StreamingLLM Xiao et al. (2023): In addition to the most recent tokens, StreamingLLM always caches the first $n$ tokens in the sequence, as they are identified as ”sink tokens”. We take $n=4$ as recommended by the paper.
5.2 End-to-End Result
Figure 3 demonstrates the comparison results of SqueezeAttention and other 3 baseline algorithms over all 7 models and 5 datasets. Full Cache (dashed line) means all tokens’ KV embeddings are fully cached during the inference, therefore, it represents a relatively good model accuracy. The blue and orange curves in each subfigure illustrate how the model accuracy changes with the KV-cache budgets ranging from 10% to 100% of the total sequence length. Note that applying different sequence-wise compression algorithms to different tasks would lead to quite different model accuracies. Therefore, for each task, we choose the best sequence-wise compression algorithm to represent the best case, and then apply SqueezeAttention on top of the best case. As shown in the figure, SqueezeAttention consistently manages to improve the model accuracies under various KV-cache budgets by reallocating the cache budgets among layers. In other words, SqueezeAttention can achieve similar inference accuracies with much less KV-cache in total.
5.3 Memory Consumption
Table LABEL:accuary_result demonstrates that in all three settings, by achieving the same model accuracy, SqueezeAttention consumes the least amount of KV-cache budgets compared with the algorithms that only compress from the sequence’s dimension. In some cases, it only takes one-third of the cache budget of H2O algorithm. Subsequently, we employ PYTORCH PROFILER to evaluate the diminished memory usage of generating one token during the inference (w/o memory usage of model weights). Figure LABEL:efficiency_result shows that SqueezeAttention can save 70% to 80% of memory usage per token compared with Full Cache method, and 25% to 66% of memory usage compared with baseline algorithms.
Table 2: Comparisons of the required KV-cache budget to achieve the best accuracy. Three models are selected to represent the small (7B), medium (20B), and large models (70B). For each task, we choose the best existing sequence-wise sparsification algorithm as the baseline.
Model Size Dataset Best Baseline Performance / Used KV Budget
Full Cache w/ SqueezeAttention w/o SqueezeAttention
@@ 5.3 Memory Consumption
5.4 Throughput of Token Generation
Since SqueezeAttention manages to save the memory cost of inference, as shown in the previous sections, we want to explore how these memory reductions can be interpreted into improvements of token throughput. We choose 2 models, i.e., Mistral-7B and Llama-70B, to represent models in small and large sizes. With a fixed content length, we increase the batch size from 1 to 224 for Mistral-7B, and 1 to 64 for Llama2. For each task, we select the best baseline algorithm, that is, Sliding Window for Mistral-7B and StreamingLLM for Llama2-70B.
Table LABEL:throughput_result shows the token throughput on 8 A100-40GB GPUs. With the same batch size, SqueezeAttention can enhance throughput by up to 2.2$\times$ for Mistral-7B and 1.4$\times$ for Llama2-70B compared to the Full Cache. Besides, SqueezeAttention also enables batch sizes up to 224 and 64 for two models, which would cause out-of-memory for the Full Cache method. The throughput comparison between best baseline and SqueezeAttention is reported in A.4.
Table 3: Generation throughput (token / s) on eight A100 GPUs of Mistral-7B and LLama2-70B with SqueezeAttention and Full Cache. To maintain the model accuracy, SqueezeAttention uses 20% of the cache budget for Mistral-7B and 30% of the cache budget for LLama2-70B. ”OOM” means out-of-memory.